OpenStudy (anonymous):

Evaluate the following expression. 6^-3

OpenStudy (anonymous):

\[\color{green}{6^{-3} = \dfrac{1}{6^3} \\ = \dfrac{1}{6 \times 6 \times 6}}\]

OpenStudy (anonymous):

ghuczek is correct.